# Bayesian mitigation of measurement errors in multi-qubit experiments

### v1

@@ VII Conclusions
In this work, we have introduced and conducted a comprehensive analysis of an efficient measurement error mitigation scheme that significantly enhances the accuracy of the readout step in multiqubit experiments on quantum computers. The algorithm combines Bayesian inference with a heuristic binary optimization cycle to construct a probability distribution for the multiqubit state populations, based on the known noise matrix affecting the measurement and state assignment. This approach effectively incorporates noise and imperfections in the post-processing of measurement outcomes. Crucially, our Bayesian framework is designed to avoid assigning unphysical values, ensuring that the mitigated outcomes remain consistent with the experimental data at each step of the mitigation algorithm. We tested the scalability and performance of the algorithm on two real superconducting qubit quantum computers, demonstrating a substantial improvement in the precision of multiqubit bitstrings with minimal computational overhead. Additionally, we have shown that it is more advantageous to apply readout mitigation techniques at the level of the detector’s analog data. For superconducting QPUs, this involves working with data in the IQ-plane rather than aggregated binary counts. We compared our method against the state-of-the-art Iterative Bayesian Unfolding (IBU) mitigation technique and found that our approach can outperform IBU, especially when incorporating analog readout data into the mitigation scheme. Although we focused on superconducting qubit devices, our mitigation approach is versatile and could be applied to other quantum computing platforms, either directly applying the mitigation procedure on the aggregated bitstring counts or using any detector analog data if available. This flexibility underscores the broad applicability and potential impact of our Bayesian measurement error mitigation framework in advancing the field of quantum computing.
@@ Acknowledgements.
We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team. This work has been partly funded by the Business Finland Co-Innovation Project 40561/31/2020 Quantum Technologies Industrial (QuTI).

### v2

@@ VI Conclusions
In this work, we have introduced and conducted a comprehensive analysis of an efficient measurement error mitigation scheme that significantly enhances the accuracy of the readout step in multiqubit experiments on quantum computers. The algorithm combines Bayesian inference with a heuristic binary optimization cycle to construct a probability distribution for the multiqubit state populations, based on the known noise matrix affecting the measurement and state assignment. Crucially, our Bayesian framework is designed to avoid assigning nonphysical values, ensuring that the mitigated outcomes remain consistent with the experimental data at each step of the mitigation algorithm. We tested the scalability and performance of the algorithm on three real superconducting qubit quantum computers, demonstrating a substantial improvement in the precision of multiqubit bitstrings at the cost of a classical computational overhead. Additionally, we have shown that it is more advantageous to apply readout mitigation techniques at the level of the detector’s analog data. For superconducting QPUs, this involves working with data in the IQ-plane rather than aggregated binary counts. We compared our method against the state-of-the-art Iterative Bayesian Unfolding (IBU) mitigation technique and the Mthree method and found that our approach can outperform IBU, when incorporating analog readout data into the mitigation scheme. Furthermore, we have shown how our method can be integrated with existing techniques to obtain even better performances. Although we focused on superconducting qubit devices, our mitigation approach is versatile and could be applied to other quantum computing platforms, either directly applying the mitigation procedure on the aggregated bitstring counts or using any detector analog data if available. This flexibility underscores the broad applicability and potential impact of our Bayesian measurement error mitigation framework in advancing the field of quantum computing.
@@ Acknowledgements.
We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team. This work has been partly funded by the Business Finland Co-Innovation Project 40561/31/2020 Quantum Technologies Industrial (QuTI). This research was partially supported by the PNRR MUR project PE0000023-NQSTI through the secondary projects “ThAnQ” and “QuCADD”.